OpenStudy (johnnydicamillo):

The theory of relativity predicts that an object whose mass is m0 when it is at rest will appear heavier when it is moving at speeds near the speed of light. When the object is moving at speed v, its mass m is given by

OpenStudy (johnnydicamillo):

The theory of relativity predicts that an object whose mass is \[m _{0}\] when it is at rest will appear heavier when it is moving at speeds near the speed of light. When the object is moving at speed v, its mass m is given by \[m=\frac{ m _{0} }{ \sqrt{1-(v^2/c^2)} }\] where c is the speed of light. Find dm/dv and explain in terms of physics what this quantity tells you

OpenStudy (johnnydicamillo):

am I just taking the derivative of this?

OpenStudy (briensmarandache):

yeah you would take the derivative

OpenStudy (johnnydicamillo):

m = m_0 - (1-(v^2/c^2))^1/2

OpenStudy (johnnydicamillo):

\[m = m_0 - (1-(v^2/c^2))^-1/2\]

OpenStudy (johnnydicamillo):

then \[m' = 1 +1/2(1-(v^2/c^2))^{-3/2}\]

OpenStudy (johnweldon1993):

Where did the minus sign come from?

OpenStudy (johnnydicamillo):

because it was originally division

OpenStudy (johnnydicamillo):

should I use the quotient rule?

OpenStudy (johnweldon1993):

\[\large \frac{m_0}{\sqrt{1 - \frac{v^2}{c^2}}} = m_0((1 - \frac{v^2}{c^2})^{-1/2})\] proceed with the chain rule *or yes if you prefer, the quotient rule

OpenStudy (johnweldon1993):

You can always write division as multiplication....but there would be no minus sign added on

OpenStudy (johnnydicamillo):

\[m_0((1-\frac{ v^2 }{ c^2 })^{-1/2}*\frac{ -1 }{ 2 }(1-\frac{ v^2 }{ c^2 }*(c^2)(2v)-v^2(2c))\]

OpenStudy (johnnydicamillo):

how wrong am I?